The length of a rectangle is 7 less than the width. The area is 30. Find the dimensions, then find the perimeter.
Answer by ikleyn(52778)

.

ANSWER.  The dimensions are  3 units (the length) and 10 units (the width).  The perimeter is  3 + 10 + 3 + 10 = 26 units.



Solution


    x*(x+7) = 30

    x^2 + 7x - 30 = 0

    (x-3)*(x+10)  = 0


The roots are 3 and -10, and only positive root works.


It is 3 units, the width.

Solved.

The solution can be guessed mentally in 2 seconds.

------------------

If you want to see other similar solved problems with detailed explanations, look into the lesson
    - Problems on the area and the dimensions of a rectangle surrounded by a strip
in this site.